An AT&T bond has 10 years until maturity, a $1,000 nominal value, a coupon rate of 8% and sells for $1,100 (coupon is paid annually).

What is the yield to maturity?
1
Expert's answer
2020-01-08T09:33:14-0500

"YTM=(C+(F-P)\/n)\/((F+P)\/n)"

Where

C - Coupon payment

F -face value

P - Price

n - years to maturity

"Coupon Rate = C\/F"

"C=0.08\u00d71000=80"

"YTM=(80+(1000-1100)\/10)\/((1000+1100)\/10=70\/210=1\/3"